Mal87:

What is the midpoint between the points (1, 3) and (5, 7)? A. (4, 4) B. (7, 9) C. (3, 5) D. (2, 2)

Hero:

Hint: Use The Midpoint Formula

Hero:

\(\left(\dfrac{x_1 + x_2}{2}\right), \left(\dfrac{y_1 + y_2}{2}\right)\)

Mal87:

I got C

Hero:

Correct

Mal87:

Thanks :)